OpenStudy (anonymous):

Calculate the equation of the line between the points (−6, −1) and (2, −1).

OpenStudy (triciaal):

Calculate the equation of the line? I can find the length of the line and the slope of the line also write the equation of the line but don't know what you mean by "calculate the equation of the line."

OpenStudy (anonymous):

To do this we use the two point formula. You substitute the values in from the two points and then simplify down to calculate and equation. \[y-y_1=\frac{y_2-y_1}{x_2-x_1}(x-x_1)\] \[y+1=\frac{-1+1}{2+6}(x+6)\] \[y=0\times(x+6)-1\] \[y=-1\] is the equation of the line